Use the definition of derivative to find the derivative of h(t) = √ 2x − 1.


1
Expert's answer
2021-04-12T18:14:41-0400

Definition of derivative is:


"h'(x) = \frac{h(x+p)-h(x)}{p}"

where: "p= \Delta x"


Let "y=h(x)=\sqrt{2x − 1}" , then


"y+\Delta y=\sqrt{2(x+\Delta x) − 1}\\
y+\Delta y=\sqrt{2x+2\Delta x − 1}\\
\Delta y=\sqrt{2x+2\Delta x − 1}-y\\
\Delta y=\sqrt{2x+2\Delta x − 1}-\sqrt{2x-1}\\
\frac{\Delta y}{\Delta x}= \frac{\sqrt{2x+2\Delta x − 1}-\sqrt{2x-1}}{\Delta x}"

Rationalising the numerator, we have:


"\frac{\Delta y}{\Delta x}= \frac{\sqrt{2x+2\Delta x − 1}-\sqrt{2x-1}}{\Delta x} \times \frac{\sqrt{2x+2\Delta x − 1}+\sqrt{2x-1}}{\sqrt{2x+2\Delta x − 1}+\sqrt{2x-1}}\\
\frac{\Delta y}{\Delta x}= \frac{2x+2\Delta x − 1-(2x-1)}{\Delta x(\sqrt{2x+2\Delta x − 1}+\sqrt{2x-1})} \\
\frac{\Delta y}{\Delta x}= \frac{2x+2\Delta x − 1-2x+1}{\Delta x(\sqrt{2x+2\Delta x − 1}+\sqrt{2x-1})} \\
\frac{\Delta y}{\Delta x}= \frac{2\Delta x}{\Delta x(\sqrt{2x+2\Delta x − 1}+\sqrt{2x-1})} \\
\frac{\Delta y}{\Delta x}= \frac{2}{\sqrt{2x+2\Delta x − 1}+\sqrt{2x-1}} \\"

Taking limits of both sides


"\lim_{\Delta x \rightarrow 0}\frac{\Delta y}{\Delta x}= \lim_{\Delta x \rightarrow 0}\frac{2}{\sqrt{2x+2\Delta x − 1}+\sqrt{2x-1}} \\
\frac{dy}{dx} =\frac{2}{\sqrt{2x − 1}+\sqrt{2x-1}} \\
\frac{dy}{dx} =\frac{2}{2\sqrt{2x-1}} \\
\frac{dy}{dx} =\frac{1}{\sqrt{2x-1}} \\"

Thus:


"h'(x) = \frac{dy}{dx} =\frac{1}{\sqrt{2x-1}} \\"